a kicker on the football team scores 1 point for making and extra point and 3 points for making a field goal. The kicker makes a total of 53 kicks for the season (Including extra points and field goals) and scores a total of 89 points. Write and solve a system of equations to find the number of extra points (x) and the number of field goals (y).
Found 2 solutions by ewatrrr, greenestamps:
Answer by ewatrrr(24785)



Hi
Question States: 53 kicks and 89 points
let x be the number of FG
 1(53-x) + 3x = 89
 2x = 36
  x = 18 FG  and 35 Kicks

35 + 3*18 = 89  checks.
Wish You the Best in your Studies.


Answer by greenestamps(13334)



The other tutor solved the problem using a single equation -- not a system of equations, as requested.

To be sure, solving the problem using a single variable is less work... but it's not what the post asked for.
x+y=53   (1) The total number of kicks was 53
3x+1y=89   (2) The point total was 89

With both equations in the form Ax+By=C, I would always use elimination instead of substitution.
      3x+y = 89
    -  x+y = 53
    ------------
      2x   = 36

I will assume you can finish from there.